OpenStudy (anonymous):

put in slope intercept form what passes through (0,7) and is parallel to y=4x-19

OpenStudy (anonymous):

Ok this is easy so: The line is parallel to y=4x -19 Parallel means that the slope of the lines must be the same Remember: In an equation of y=mx+b m is the slope and b is the y-intercept for your equation m=4 because y=(4)x-19 So so far we know that the equation for your new line is y=4x+b Now to find "b" we need to know an x and y coordinate but wait they gave us one so plug those in for x and y to solve for b: so: 7=4(0)+b 7=0+b 7=b So we know b=7 So our equation in form y=mx+b where m is the slope is: y=4x+7 ANSWER: y=4x+7